How many milliliters of 10.0 M HNO3 solution must be added to a 1.0 L buffer that is .01 M acetic acid ( HC2H3O2) and .2M sodium acetate (NaC2H3O2) to bring the pH to 5.6? the pKa of acetic acid is 4.75
Answer #1

pH of the solution after addition of HNO3 = 5.6 Let x be the number moles of HNO3 to be added to the buffer to change the pH.

Number of moles of CH3COOH = M*V = 0.01 M * 1.0 L = 0.01 mol Number of moles of CH3COO-= M*V = 0.2 M * 1.0 L = 0.2 mol CH3COO